A hemispherical dome holds a volume of 486 pie ft^3. Find the surface area. (omit the area of the circle)
* please help me!!!***

Answer by mathstutor494(120)

A hemispherical dome holds a volume of 486 pie ft^3. Find the surface area.
Volume of Semisphere = 2/3 pi r^3
> 486 = 2/3 pi r^3
> r^3 = (486*3)/(2*pi)
Hence r = 9/ 1.46 > 6.16
Surface Area of a Semisphere = 2 pi r^2
Area of a circle = pi r^2
So Surface Area of a Semisphere less Area of a circle
> 2 pi r^2 - pi r^2
> pi r^2
Substituting values of pi & r from above
> 3.14* (6.16)^2
> 119.15 ft^2